# MUSES: 3D-Controllable Image Generation via Multi-Modal Agent Collaboration

### v2

@@ Abstract
Despite recent advancements in text-to-image generation, most existing methods struggle to create images with multiple objects and complex spatial relationships in 3D world. To tackle this limitation, we introduce a generic AI system, namely MUSES, for 3D-controllable image generation from user queries. Specifically, our MUSES addresses this challenging task by developing a progressive workflow with three key components, including (1) Layout Manager for 2D-to-3D layout lifting, (2) Model Engineer for 3D object acquisition and calibration, (3) Image Artist for 3D-to-2D image rendering. By mimicking the collaboration of human professionals, this multi-modal agent pipeline facilitates the effective and automatic creation of images with 3D-controllable objects, through an explainable integration of top-down planning and bottom-up generation. Additionally, we find that existing benchmarks lack detailed descriptions of complex 3D spatial relationships of multiple objects. To fill this gap, we further construct a new benchmark of T2I-3DisBench (3D image scene), which describes diverse 3D image scenes with 50 detailed prompts. Extensive experiments show the state-of-the-art performance of MUSES on both T2I-CompBench and T2I-3DisBench, outperforming recent strong competitors such as DALL-E 3 and Stable Diffusion 3. These results demonstrate a significant step of MUSES forward in bridging natural language, 2D image generation, and 3D world.
@@ Introduction
2D Layout Planning via In-Context-Learning. We start with 2D layout planning to determine the 2D location of objects. Apparently, asking the LLM directly to generate an object layout is not ideal, as it may struggle with managing bounding boxes in images. Hence, we leverage In-Context Learning (Dong et al. 2022), allowing the LLM to follow instructions with a few examples from a 2D layout shop. We use the NSR-1K dataset (Feng et al. 2024) as the 2D layout shop, since it contains over 40,000 entries with diverse prompts, objects, and the corresponding bounding boxes. First, we use the CLIP (Radford et al. 2021) text encoder to compare similarities of user input and text prompts in the shop. Then, we select the top five 2D layouts with the highest similarity scores. Finally, we feed these contextual layouts along with instructions to the LLM, guiding it to generate a 2D layout of the user input. Such a comprehensive approach results in a contextually-relevant 2D layout, forming the foundation for the subsequent 3D lifting.
@@
CompAgent (Wang et al. 2024) 0.6994 0.5740 0.6927 0.3138 - 0.3102 -
SD v1.4 (Rombach et al. 2022a) 0.3765 0.3576 0.4156 0.1350 0.3336 0.3079 0.4418
SD v2 (Rombach et al. 2022b) 0.5065 0.4221 0.4922 0.1703 0.3205 0.3096 0.4849
@@
MUSES(Ours) 0.7668 0.5762 0.7206 0.3544 0.3893 0.3214 0.5673
@@
Structured Diffusion (Feng et al. 2023) 0.2202 (0.15) 0.2308 (0.14) 0.1770 (0.08) 0.3040 (0.21) 0.1690 (0.18)
RPG-Diffusion (Yang et al. 2024b) 0.2444 (0.18) 0.3100 (0.18) 0.2275 (0.15) 0.2295 (0.25) 0.2105 (0.12)
LayoutGPT (Feng et al. 2024) 0.2957 (0.14) 0.3246 (0.14) 0.1659 (0.12) 0.4587 (0.18) 0.2337 (0.10)
Playground v2.5 (Li et al. 2024a) 0.1776 (0.19) 0.2300 (0.19) 0.1502 (0.17) 0.2050 (0.28) 0.1250 (0.15)
DALL-E 3 (Betker et al. 2023) 0.1925 (0.23) 0.1960 (0.15) 0.1720 (0.18) 0.2027 (0.29) 0.1990 (0.29)
Hunyuan-DiT (Li et al. 2024b) 0.2099 (0.22) 0.2130 (0.19) 0.1780 (0.21) 0.2509 (0.29) 0.1975 (0.18)
Midjourney v6.0 (Midjourney 2024) 0.3538 (0.35) 0.3820 (0.39) 0.3322 (0.27) 0.3880 (0.36) 0.3130 (0.38)
Stable Diffusion V3 (Esser et al. 2024) 0.4443 (0.36) 0.5840 (0.42) 0.3590 (0.26) 0.5081 (0.40) 0.3260 (0.33)
MUSES(Ours) 0.5192 (0.62) 0.6180 (0.61) 0.4610 (0.68) 0.5647 (0.62) 0.4330 (0.57)
@@
Results on T2I-CompBench 0.3893 0.3647 0.3816 0.3893 0.3536 0.3567 0.3640 0.3517 0.3439 0.3471 0.3406
Results on T2I-3DisBench 0.5192 0.3852 0.4025 0.4580 0.3837 0.4988 0.4437 0.3455 0.3558 0.3301 0.2980
@@
Implementation Details. Our MUSES is designed to be modular and extensible, allowing for the integration of various LLMs, CLIPs and ControlNets. In our experimental setup, we employed Llama-3-8B (AI@Meta 2024) for 3D layout planning, ViT-L/14 for image and text encoding, ViT-B/32 for orientation calibration, and SD 3 ControlNet (Zhang, Rao, and Agrawala 2023) for 3D controllable image generation. For Llama-3-8B, we set top_p to 0.1 and temperature to 0.2 to ensure precise, consistent, and reliable outputs. For SD 3 ControlNet, we set inference step to 20 and control scale to 0.5, as discussed in parameter ablation in Supplementary Material Section D. The Blender’s parameter conversion rules and settings are also specified in Supplementary Material Section C. Additionally, we employed Mini-InternVL 1.5 (Chen et al. 2024a) for automated evaluation on T2I-3DisBench (implementation details in Supplementary Material). All experiments are conducted on Ubuntu 20.04 system with 8 NVIDIA RTX 3090 GPUs.
@@
Ablation Studies of Different Components. As shown in Table 3, our full system achieves the best performance, and removing any component degrades the results. (1) Object Depth Planning is essential, as its removal leads to poor 3D spatial representation in Blender, affecting both datasets. (2) Object Orientation Planning is critical for prompts containing orientation information, with a score dropping on T2I-3DisBench when removed. (3) Camera View Planning also affects performance on T2I-3DisBench, since it contains camera information. (4) Retrieve-Generate Decision Tree significantly influences performance on both datasets, highlighting the importance of high-quality 3D models. (5) Face-Camera Calibration is the most important; its removal causes sharp performance drops as models lose correct orientation in the 3D layout. (6) CLIP Fine-tuning During Calibration improves CLIP’s accuracy in determining object orientations, thus enhancing the orientation accuracy of our final image. (7) Co-ablation of Multiple Components (last four columns) shows that removing the Model Engineer has the most significant impact, resulting in poor object shaping and orientation. Removing the Layout Manager also notably degrades performance. These findings demonstrate the interconnected nature of our system’s components and their effects on overall performance, with each component being crucial for 3D controllable image generation.
@@ Acknowledgments
Appendix A Full Prompts for Planning
@@ Appendix B T2I-3DisBench Benchmark
Owing to the absence of a suitable textual dataset involving multiple objects with various orientations, 3D spatial relationships, and camera views, we construct our own benchmark, namely T2I-3DisBench (3D image scene Benchmark). The benchmark construction process begins with the careful crafting of 10 such complex and detailed prompts. To expand the dataset, we leveraged Claude AI [^1] https://www.anthropic.com/claude to imitate and generate remaining 40 similar texts, which are subsequently refined by human experts to ensure quality, diversity and consistency. Fig. 7 presents some representative examples in our T2I-3DisBench textual dataset. For evaluation metrics, we find that traditional metrics such as CLIPScore (Radford et al. 2021) or BLIP-CLIP (Li et al. 2022) lacked the necessary precision to capture nuanced details like object orientations or camera views, and the metrics of T2I-compBench (Huang et al. 2023) are inadequate and inaccurate for assessing detailed 3D features. Hence, we employ InternVL (Chen et al. 2024b), a Vision Large Language Model (VLLM), to score the generated images across four dimensions: object count, object orientation, 3D spatial relationship, and camera view. We feed InternVL with the following instruction:
@@
How well does the image match the text? You need to consider (1) object count, (2) object orientation, (3) 3D spatial relationship between objects, (4) camera view. Return a tuple (”score”, X.XXXX), with a float number between 0 and 1. The higher value represents higher text-image alignment.
@@ Appendix C Implementation details in Blender
After setting all the parameters, we now successfully assemble the 3D scene. We are able to render the complete 3D scene into a 2D image with accuracy using the $bpy.ops.render.render(True)$ function. This 2D rendering is further transformed into both a depth map and an edge map, which are used for fine-grained control in the ControlNet (Zhang, Rao, and Agrawala 2023). To generate the depth map, we use a depth node within Blender to capture the Z-depth values of the 3D objects. This depth information is crucial for understanding the spatial relationships between objects in the scene. For the edge map, we employ OpenCV to detect the contours in the rendered image. This edge map highlights the boundaries and shapes of objects, providing additional information that aids in the precise control and manipulation of the 3D scene. Our specific codes are presented in the following.
@@ Appendix D Ablation of ControlNet Parameters
We conduct comparative experiments with different control scales and inference steps of ControlNets to determine the optimal parameter settings. The $controlscale$ parameter ranges from 0.1 to 0.9, and the $inferencesteps$ parameter ranges from 5 to 25. As shown in Fig. 8, we select several representative parameter values and visualize the comparison results. The results indicate that increasing the control scale enhances the alignment of the generated images with the input condition images. However, this improvement comes at the cost of reduced image quality. To balance these effects, the optimal control scale parameter is identified as 0.5. For inference steps, the quality of the generated images improves with an increasing number of steps. However, once the number of steps exceeds 20, the image quality gains plateau. Therefore, 20 inference steps are chosen as the optimal inference parameter in our experimental setup.
@@ Appendix E More Qualitative Comparisons
Figure 9 and Figure 10 present more qualitative comparisons between our MUSES and existing state-of-the-art methods, including both open-source models and commercial API products, such as Stable Diffusion V3 (Esser et al. 2024), DALL-E 3 (Betker et al. 2023), and Midjourney v6.0 (Midjourney 2024). Obviously, our systematic collaborative approach generates images that are more faithful to the details of the text, in terms of object count, orientation, 3D spatial relationships, and camera view. Additionally, when dealing with such complex and detailed prompts, our system incorporates a human verification step following the 2D-to-3D Layout Manager’s initial planning. This human-in-the-loop approach further ensures that the planned layout reasonably represents the details in the user input. It also demonstrates the potential for human-AI collaboration in tackling such complex visual generation tasks.

### v3

@@ Abstract
Despite recent advancements in text-to-image generation, most existing methods struggle to create images with multiple objects and complex spatial relationships in 3D world. To tackle this limitation, we introduce a generic AI system, namely MUSES, for 3D-controllable image generation from user queries. Specifically, our MUSES addresses this challenging task by developing a progressive workflow with three key components, including (1) Layout Manager for 2D-to-3D layout lifting, (2) Model Engineer for 3D object acquisition and calibration, (3) Image Artist for 3D-to-2D image rendering. By mimicking the collaboration of human professionals, this multi-modal agent pipeline facilitates the effective and automatic creation of images with 3D-controllable objects, through an explainable integration of top-down planning and bottom-up generation. Additionally, we find that existing benchmarks lack detailed descriptions of complex 3D spatial relationships of multiple objects. To fill this gap, we further construct a new benchmark of T2I-3DisBench (3D image scene), which describes diverse 3D image scenes with 50 detailed prompts. Extensive experiments show the state-of-the-art performance of MUSES on both T2I-CompBench and T2I-3DisBench, outperforming recent strong competitors such as DALL-E 3 and Stable Diffusion 3. These results demonstrate a significant step of MUSES forward in bridging natural language, 2D image generation, and 3D world. Our codes and models will be released soon.
@@ Introduction
2D Layout Planning via In-Context-Learning. We start with 2D layout planning to determine the 2D location of objects. Apparently, asking the LLM directly to generate an object layout is not ideal, as it may struggle with managing bounding boxes in images. Hence, we leverage In-Context Learning (Dong et al. 2022), allowing the LLM to follow instructions with a few examples from a 2D layout shop. We use the NSR-1K dataset (Feng et al. 2024) as the 2D layout shop, since it contains over 40,000 entries with diverse prompts, objects, and the corresponding bounding boxes. However, the original NSR-1K dataset lacks layouts in 3D scenes and complex scenes, so we manually add some prompts with corresponding layouts (details in Supplementary Material). First, we use the CLIP (Radford et al. 2021) text encoder to compare similarities of user input and text prompts in the shop. Then, we select the top five 2D layouts with the highest similarity scores. Finally, we feed these contextual layouts along with instructions to the LLM, guiding it to generate a 2D layout of the user input. Such a comprehensive approach results in a contextually-relevant 2D layout, forming the foundation for the subsequent 3D lifting.
@@
CompAgent (Wang et al. 2024) 0.7400 0.6305 0.7102 0.3698 - 0.3104 -
@@
DALL-E 3 (Betker et al. 2023) 0.7785 0.6205 0.7036 0.2865 - 0.3003 -
SD v3 (Esser et al. 2024) 0.8085 0.5793 0.7317 0.3144 0.4026 0.3131 0.6088
MUSES (Ours) 0.8726 0.6812 0.8081 0.4756 0.4639 0.3226 0.7720
@@
Structured Diffusion (Feng et al. 2023) 0.1862 (0.15) 0.2160 (0.14) 0.1647 (0.08) 0.1773 (0.21) 0.1866 (0.18)
RPG-Diffusion (Yang et al. 2024b) 0.2753 (0.18) 0.3209 (0.18) 0.2533 (0.15) 0.3195 (0.25) 0.2075 (0.12)
LayoutGPT (Feng et al. 2024) 0.2348 (0.14) 0.2937 (0.14) 0.2058 (0.12) 0.2783 (0.18) 0.1613 (0.10)
Playground v2.5 (Li et al. 2024a) 0.2344 (0.19) 0.3587 (0.19) 0.1887 (0.17) 0.2071 (0.28) 0.1830 (0.15)
DALL-E 3 (Betker et al. 2023) 0.2627 (0.23) 0.3013 (0.15) 0.2363 (0.18) 0.2370 (0.29) 0.2757 (0.29)
Hunyuan-DiT (Li et al. 2024b) 0.2780 (0.22) 0.3496 (0.19) 0.2517 (0.21) 0.2598 (0.29) 0.2510 (0.18)
Midjourney v6.0 (Midjourney 2024) 0.3760 (0.35) 0.4470 (0.39) 0.3438 (0.27) 0.3518 (0.36) 0.3613 (0.38)
SD v3 (Esser et al. 2024) 0.4206 (0.36) 0.5383 (0.42) 0.3303 (0.26) 0.4600 (0.40) 0.3537 (0.33)
MUSES (Ours) 0.6156 (0.62) 0.7488 (0.61) 0.4709 (0.68) 0.7207 (0.62) 0.5220 (0.57)
@@
Multiple Control Scales
3D Layout Shop Expansion
Results on T2I-CompBench 0.4639 0.4127 0.4609 0.4636 0.4234 0.3828 0.4289 0.3893 0.3925 0.3641 0.3427 0.3093
Results on T2I-3DisBench 0.6156 0.4981 0.4938 0.5245 0.5360 0.4019 0.5026 0.5192 0.4405 0.3748 0.3661 0.3178
@@
Implementation Details. Our MUSES is designed to be modular and extensible, allowing for the integration of various LLMs, CLIPs and ControlNets. In our experimental setup, we employed Llama-3-8B (AI@Meta 2024) for 3D layout planning, ViT-L/14 for image and text encoding, ViT-B/32 for orientation calibration, and SD 3 ControlNet (Zhang, Rao, and Agrawala 2023) for controllable image generation. For Llama-3-8B, we set top_p to 0.1 and temperature to 0.2 to ensure precise, consistent, and reliable outputs. For SD 3 ControlNet, we set inference steps to 20 and control scales from 0.5 to 0.9, as discussed in parameter ablation in Supplementary Material Section D. During evaluation, we select the best one from the five images conditioned on different control scales using benchmark metrics (e.g., T2I-CompBench and T2I-3DisBench). The Blender’s parameter conversion rules and settings are also specified in Supplementary Material Section C. Additionally, we employed Mini-InternVL 1.5 (Chen et al. 2024a) for automated evaluation on T2I-3DisBench (details in Supplementary Material). All experiments are conducted on Ubuntu 20.04 system equipped with 8 NVIDIA RTX 3090 GPUs.
@@
Ablation Studies. As shown in Table 3, our full system achieves the best performance. (1) Object Depth Planning is essential, as its removal leads to poor 3D spatial representation in Blender, affecting both datasets. (2) Object Orientation Planning is critical for prompts containing orientation information, with a score dropping on T2I-3DisBench when removed. (3) Camera View Planning affects performance mostly on T2I-3DisBench, since it contains camera information. (4) Retrieve-Generate Decision Tree significantly influences performance on both datasets, highlighting the importance of high-quality 3D models. (5) Face-Camera Calibration is the most important; its removal causes sharp performance drops as 3D models lose correct orientation specified in the 3D layout. (6) CLIP Fine-tuning During Calibration improves CLIP’s accuracy in determining object orientations, thus enhancing the orientation accuracy of our final image. (7) Multiple Control Scales effectively improves performance on both benchmarks. (8) 3D Layout Shop Expansion helps the LLM generate 3D and complex layouts better; it obviously works. (9) Co-ablation of Multiple Components (last three columns) shows that removing the Model Engineer has the most significant impact, resulting in poor object shaping and orientation. Removing the Layout Manager also notably degrades performance. Removing all components will result in the lowest score. These findings demonstrate that each component is crucial for our 3D controllable image generation system.
@@ Acknowledgments
Appendix A Full Prompts for 3D Layout Planning
@@ Appendix B T2I-3DisBench Benchmark
Owing to the absence of a suitable textual dataset involving multiple objects with various orientations, 3D spatial relationships, and camera views, we construct our own benchmark, namely T2I-3DisBench (3D image scene Benchmark). The benchmark construction process begins with the careful crafting of 10 such complex and detailed prompts. To expand the dataset, we leverage the Claude AI as well to imitate and generate remaining 40 similar texts, which are subsequently refined by human experts to ensure quality, diversity and consistency. Fig. 8 presents some representative examples in our T2I-3DisBench textual dataset. For evaluation metrics, we find that traditional metrics such as CLIPScore (Radford et al. 2021) or BLIP-CLIP (Li et al. 2022) lack the necessary precision to capture nuanced details like object orientations or camera views, and the metrics of T2I-compBench (Huang et al. 2023) are inadequate and inaccurate for assessing detailed 3D features. Hence, we employ InternVL (Chen et al. 2024b), a Vision Large Language Model (VLLM), to score the generated images across four dimensions: object count, object orientation, 3D spatial relationship, and camera view. Specifically, we feed the InternVL with the following instructions:
@@
How well does the image match the text? You need to consider (1) object count, (2) object orientation, (3) 3D spatial relationship between objects, (4) camera view. Return a tuple (”score”, X.XXXX), with the float number between 0 and 1, and higher scores representing higher text-image alignment.
@@ Appendix C Implementation details in Blender
After setting all the parameters, we now successfully assemble the 3D scene. We are able to render the complete 3D scene into a 2D image with accuracy using the $bpy.ops.render.render(True)$ function. This 2D rendering is further transformed into both a depth map and an edge map, which are used for fine-grained control in the ControlNet (Zhang, Rao, and Agrawala 2023). To generate the depth map, we use a depth node within Blender to capture the Z-depth values of the 3D objects. This depth information is crucial for understanding the spatial relationships between objects in the scene. For the edge map, we employ OpenCV to detect the contours in the rendered image. This edge map highlights the boundaries and shapes of objects, providing additional information that aids in the precise manipulation of the 3D scene. Codes are presented in the following.
@@ Appendix D Ablation of ControlNet Parameters
We conduct comparative experiments with different control scales and inference steps of ControlNets to determine the optimal parameter settings. The $control\ scale$ parameter ranges from 0.1 to 1.0, and the $inference\ steps$ parameter ranges from 5 to 30. As shown in Fig. 9, we select several representative parameter values and visualize the comparison results. The results indicate that increasing the control scale enhances the alignment of the generated images with the input condition images. When the control scale reaches 0.5, the details of the generated image can already be well controlled. And it is difficult to determine which control scale results in images with better quality after 0.5. Therefore, during the experiments, we set the control scale of SD 3 ControlNet in the interval of [0.5, 0.9]. That is, each run generates five images with different control scales, and then evaluates them using benchmark (e.g., T2I-CompBench and T2I-3DisBench) metrics to select the best image with the highest score. As for inference steps, the quality of the generated images improves with an increasing number of steps. However, once the number of steps exceeds 20, the image quality gains plateau. Therefore, 20 inference steps are chosen as the optimal inference steps in our experimental setup.
@@ Appendix E More Qualitative Comparisons
Figure 10 and Figure 11 present more qualitative comparisons between our MUSES and existing state-of-the-art methods, including both open-source models and commercial API products, such as Stable Diffusion V3 (Esser et al. 2024), DALL-E 3 (Betker et al. 2023), and Midjourney v6.0 (Midjourney 2024). Obviously, our systematic collaborative approach generates images that are more faithful to the details of the text, in terms of object count, orientation, 3D spatial relationships, and camera view. Even the state-of-the-art methods fail to generate precise images that accurately represent the complex 3D details in the input text prompts.